# Adaptive Teaming in Multi-Drone Pursuit: Simulation, Training, and Deployment

### v1

@@ 5 Conclusion
In this paper, we define and formalize the adaptive teaming in the multi-drone pursuit problem and introduce AT-MDP framework, a comprehensive framework that integrates simulation, training, and real-world deployment to address this challenge. AT-MDP framework provides a user-friendly and flexible environment configurator, a distributed training framework incorporating an extensive algorithm zoo and an unseen drone zoo, and a real-world deployment system leveraging edge computing and Crazyflie drones. To the best of our knowledge, AT-MDP framework is the first adaptive teaming framework designed for continuous-action multi-drone pursuit tasks, extending multi-agent coordination beyond pre-coordinated strategies. Our experiments in four environments of increasing difficulty validate the effectiveness of our proposed baseline methods, demonstrating their adaptability to unseen teammates and superior performance compared to existing approaches. Furthermore, real-world experiments confirm the feasibility of AT-MDP framework in physical drone systems, bridging the gap between simulation and real-world deployment.
Limitations and Future Work. While AT-MDP framework successfully bridges simulation and real-world deployment, the current real-world system is relatively simple, relying on Crazyflie drones without onboard perception, LiDAR, or computer vision capabilities. Furthermore, the scenarios tested, while diverse, could be further enhanced to better reflect real-world complexities. Future work will focus on integrating perception modules, expanding the framework to more realistic environments, and developing a multi-drone system equipped with edge computing devices, such as Jetson, to enable fully autonomous adaptive teaming.
6 Impact Statement
This paper advances the field of machine learning and multi-robot systems by introducing Adaptive Teaming in Multi-Drone Pursuit (AT-MDP) and proposing AT-MDP framework, a comprehensive framework for simulation, training, and real-world deployment. Our work contributes to multi-drone system by enabling multiple drones to dynamically coordinate with unseen teammates in continuous-action real-world tasks.
The potential societal benefits of this research include improvements in search-and-rescue operations, disaster response, border surveillance, and autonomous security systems, where adaptive coordination among unseen agents can enhance efficiency and safety. Additionally, the open-source implementation of our framework fosters reproducibility and further research in adaptive teaming and multi-agent collaboration.
From an ethical perspective, while multi-drone pursuit has positive real-world applications, the technology also has potential risks if misused, such as surveillance or adversarial applications. We emphasize the importance of responsible AI development and adherence to ethical guidelines in deploying autonomous systems. Future work should consider human-in-the-loop oversight and safety constraints to ensure the ethical use of AI-driven multi-agent systems.
Appendix A Related Work
Table 4: Comparison of AT-MDP framework with related works. Grey rows represent literature related to multi-drone pursuit, while pink rows highlight adaptive teaming studies from the machine learning field. “AT w/o TM” denotes adaptive teaming without teammate modelling, while “AT w/ TM” refers to adaptive teaming with teammate modelling.
Related Work Problem Setting Task Method
# Learner # Unseen # Evader Action Space Main Related Task Real-world? AT w/o TM? AT w/ TM?
Voronoi Partitions (Zhou et al., 2016) Multi 0 1 Continuous Pursuit–evasion Game No No No
Bio-pursuit (Janosov et al., 2017b) Multi 0 Multi Continuous Prey–predator Game No No No
Uncertainty-pursuit (Shah & Schwager, 2019) Multi 0 1 Continuous Pursuit–evasion Game No No No
M3DDPG (Li et al., 2019) Multi 0 1 Continuous Prey–predator Game No No No
Pursuit-TD3(de Souza et al., 2021) Multi 0 1 Continuous Multi-drone Pursuit Yes No No
DACOOP-A(Zhang et al., 2023) Multi 0 1 Discrete Multi-drone Pursuit Yes No No
GM-TD3 (Zhang et al., 2024) Multi 0 1 Continuous Prey–predator Game No No No
DualCL (Chen et al., 2024) Multi 0 1 Continuous Multi-drone Pursuit No No No
HOLA-Drone (Li et al., 2024a) 1 Multi Multi Continuous Multi-drone Pursuit Yes Yes No
Other-play (Hu et al., 2020) 1 1 0 Discrete Lever Game; Hanabi No Yes No
Overcooked-AI  (Carroll et al., 2019) 1 1 0 Discrete Overcooked No Yes No
TrajDi (Lupu et al., 2021) 1 1 0 Discrete Overcooked No Yes No
MEP (Zhao et al., 2023) 1 1 0 Discrete Overcooked No Yes No
LIPO (Charakorn et al., 2023) 1 1 0 Discrete Overcooked No Yes No
COLE (Li et al., 2024b) 1 1 0 Discrete Overcooked No Yes No
ZSC-Eval (Wang et al., 2024d) 1 1 0 Discrete Overcooked No Yes No
PLASTIC (Barrett et al., 2017) 1 Multi Multi Discrete Prey-predator Game No No Yes
AATeam (Chen et al., 2020) 1 1 2 Discrete Half Field Offense No No Yes
LIAM (Papoudakis et al., 2021) 1 Multi Multi Discrete LBF; Prey-predator Game No No Yes
GPL (Rahman et al., 2021) 1 Multi Multi Discrete LBF; Wolfpack; FortAttack No No Yes
CIAO (Wang et al., 2024b) 1 Multi Multi Discrete LBF; Wolfpack No No Yes
NAHT (Wang et al., 2024a) Multi Multi Multi Discrete StarCraft2 No No Yes
Our ATMDP Multi Multi Multi Continuous Multi-drone Pursuit Yes Yes Yes
In this work, we provide a comprehensive review of related research on multi-drone pursuit and adaptive teaming in machine learning, with a detailed comparison presented in Table 1.
Multi-agent pursuit-evasion. Multi-agent pursuit-evasion is closely related to the multi-drone pursuit task. Most existing methods rely on pre-coordinated strategies specifically designed for particular pursuit-evasion scenarios. Traditional approaches often rely on heuristic (Janosov et al., 2017b) or optimisation-based strategies (Zhou et al., 2016; Shah & Schwager, 2019). For example, Janosov et al. (2017b) proposes a bio-inspired model that uses local interaction rules to enhance group chasing success in Prey–Predator Games. Similarly, the Voronoi partitions algorithm (Zhou et al., 2016) and the uncertainty-pursuit algorithm (Shah & Schwager, 2019) employ decentralised frameworks to optimise the evader’s Voronoi partition and reachable area, respectively. In recent years, deep reinforcement learning (DRL) has been widely adopted for pre-coordinated multi-drone pursuit tasks. M3DDPG (Li et al., 2019) and GM-TD3 (Zhang et al., 2024) extend standard DRL algorithms, such as TD3 (Fujimoto et al., 2018) and DDPG (Lillicrap, 2015), specifically for multi-agent pursuit in simulated environments. Pursuit-TD3 (de Souza et al., 2021) applies the TD3 algorithm to pursue a target with multiple homogeneous agents, validated through both simulations and real-world drone demonstrations. Zhang et al. (2023) introduces DACOOP-A, a cooperative pursuit algorithm that enhances reinforcement learning with artificial potential fields and attention mechanisms, validated in real-world drone systems. DualCL (Chen et al., 2024) addresses multi-UAV pursuit-evasion in diverse environments and demonstrates zero-shot transfer capabilities to unseen scenarios, though only in simulation. The most recent work, HOLA-Drone (Li et al., 2024a), claims to be the first zero-shot coordination framework for multi-drone pursuit. However, it is limited to controlling a single learner, restricting its applicability to broader multi-agent settings.
Adaptive Teaming. The adaptive teaming paradigm can be broadly categorised into two approaches: adaptive teaming without teammate modelling (AT w/o TM) and adaptive teaming with teammate modelling (AT w/ TM), which correspond to the zero-shot coordination (ZSC) and ad-hoc teamwork (AHT) problems in the machine learning community, respectively. AT w/o TM focuses on enabling agents to coordinate with unseen teammates without explicitly modelling their behaviours. Other-Play (Hu et al., 2020) introduces an approach that leverages symmetries in the environment to train robust coordination policies, applied to discrete-action tasks like the Lever Game and Hanabi. Similarly, methods such as Overcooked-AI (Carroll et al., 2019), TrajDi (Lupu et al., 2021), MEP (Zhao et al., 2023), LIPO (Charakorn et al., 2023), and ZSC-Eval (Wang et al., 2024d) study collaborative behaviours in Overcooked, where agents learn generalisable coordination strategies with diverse unseen partners. While these approaches demonstrate promising results, they are limited to single-learner frameworks in simplified, discrete-action domains like Overcooked and Hanabi. They lack scalability to multi-agent settings, continuous action spaces, and the complexities of real-world applications.
AT w/ TM, on the other hand, explicitly models the behaviour of unseen teammates to facilitate effective collaboration. Early methods like PLASTIC (Barrett et al., 2017) reuse knowledge from previous teammates or expert input to adapt to new teammates efficiently. AaTeam (Chen et al., 2020) introduces attention-based neural networks to dynamically process and respond to teammates’ behaviours in real-time. More advanced approaches, such as LIAM (Papoudakis et al., 2021), employ encoder-decoder architectures to model teammates using local information from the controlled agent. GPL (Rahman et al., 2021) and CIAO (Wang et al., 2024b) leverage graph neural networks (GNNs) to address the challenges of dynamic team sizes in AHT. Extending this further, NAHT (Wang et al., 2024a) enables multiple learners to collaborate and interact with diverse unseen partners in N-agent scenarios. Despite their progress, these methods remain confined to discrete action spaces and simulated benchmarks, limiting their applicability to real-world, continuous-action tasks.
Figure 4: Overview of our proposed open-ended population training algorithm.
Appendix B Open-ended Population Training Algorithm
Building on the concept of preference hypergraphs (Li et al., 2024a), we use the preference hypergraph to represent the population and assess the coordination ability of each node. The preference hypergraph $\mathcal{P}\mathcal{G}$ is derived from the hypergraph $\mathcal{G}$, where each node has a direct outgoing hyperedge pointing to the teammates with whom it achieves the highest weight in $\mathcal{G}$. Formally, $\mathcal{P}\mathcal{G}$ is defined by the tuple $(\Pi,\mathcal{E}_{\mathcal{P}})$, where the node set $\Pi$ represents the strategies, and $\mathcal{E}_{\mathcal{P}}$ denotes the set of outgoing hyperedges. As shown in the right subfigure of Fig. 5, the dotted line highlights the outgoing edge. For instance, node $2$ has a single outgoing edge $(2,3,5,4)$ because it achieves the highest outcome, i.e., a weight of 45, with those teammates in $\mathcal{G}$, as depicted in the left subfigure.
Intuitively, a node in $\mathcal{P}\mathcal{G}$ with higher cooperative ability will have more incoming hyperedges, as other agents prefer collaborating with it to achieve the highest outcomes. Therefore, we extend the concept of preference centrality (Li et al., 2023) to quantify the cooperative ability of each node. Specifically, for any node $i\in\Pi$, the preference centrality is defined as
$\eta_{\Pi}(i)=\frac{d_{\mathcal{P}\mathcal{G}}(i)}{d_{\mathcal{G}}(i)},$ (1)
Definition B.1 (Preference Optimal).
The max-min preference oracle updates the learner set by solving:
where the objective $\phi_{\mathcal{M}}(\cdot)$ is derived using the extended curry operator (Balduzzi et al., 2019), originally designed for two-player games, and is expressed as:
$\mathcal{N}^{*}=\arg\max_{\mathcal{N}}\phi(\mathcal{N},\mathcal{M}^{*}).$
Open-Ended Population Training algorithm. To achieve robust adaptability and dynamic coordination in multi-agent systems, we integrate the max-min preference oracle into an open-ended learning framework, referred to as the Open-ended Population Training (OPT) algorithm. The OPT algorithm dynamically adjusts the training objective as the population evolves, enabling continuous improvement and effective coordination with unseen partners. Unlike conventional fixed-objective training, the OPT approach iteratively expands the strategy population $\Pi$ and refines the learner set $\mathcal{N}$. At each generation $t$, the framework recalibrates the training objective $\phi$ based on new extended population $\Pi_{t}$ to account for the evolving interactions among agents within the population.
As shown in Fig. 5, the Open-Ended Population Training (OPT) algorithm consists of two key modules: the min-step solver and the max-step trainer. At each generation $t$, the updated learner set $\mathcal{N}_{t}$ from the previous generation $t-1$ is incorporated into the population $\Pi_{t-1}$, resulting in an expanded population $\Pi_{t}$.
Figure 6: Illustration of four multi-drone pursuit environments in real world. The environments vary in the number of pursuers (p), evaders (e), and obstacles (o), denoted as 4p2e3o, 4p2e1o, 4p2e5o, and 4p3e5o. Each setup introduces different levels of complexity, testing the adaptability and coordination capabilities of the agents.
Appendix C Unseen Drone Zoo
@@
Traditional Method: VICSEK Drone. Based on the commonly used VICSEK algorithm (Janosov et al., 2017a; Zhang et al., 2023; Li et al., 2024a), the VICSEK Drone adopts a bio-inspired approach to mimic swarm-like behaviours. It computes and updates a velocity vector directed towards the evader, optimising the tracking path based on the agent’s current environmental state. To avoid nearby obstacles or agents, the VICSEK Drone applies repulsive forces with varying magnitudes. While the calculated velocity vector includes both magnitude and orientation, only the orientation is implemented in our experiments, making it a scalable and practical teammate model for multi-drone coordination.
@@
Appendix D Environment Configurator
Appendix E Real-world Deployment
To validate the feasibility of our algorithms in real-world scenarios, we design and conduct hardware experiments using Crazyflie drones, the FZMotion motion capture system, and the Crazyswarm validation platform. The FZMotion motion capture system is responsible for real-time position measurement of the drones. It transmits the positional data in point cloud format to Crazyswarm, where the information is processed as input for the algorithm. The Crazyswarm platform is deployed on two different edge nodes, i.e. a laptop ( Lenovo ThinkPad T590) and a Jetson Orin Nano, which handles the reception of drone position data from the motion capture system, executes the adaptive teaming algorithm, and transmits control commands to the Crazyflie drones via Crazyradio PA. Upon receiving these commands, the Crazyflie drones execute the prescribed maneuvers. The onboard Mellinger controller ensures accurate trajectory tracking, allowing the drones to follow the control signals with high precision. This real-world deployment setup enables the direct evaluation of our learned policies in physical drone systems, bridging the gap between simulation and real-world execution.
Table 5: Implementation hyperparameters of ATM baseline algorithm.
@@ Appendix E Real-world Deployment
Appendix F Implementation Details of Adaptive Teaming with Modeling
Adaptive Teaming with Modeling (ATM) extends the MAPPO algorithm (Yu et al., 2022) by incorporating an additional teammate modeling network $f$. This network generates team encoding vectors to represent the behavioral characteristics of unseen teammates, improving coordination in multi-drone pursuit. The modeling network $f$ processes three types of inputs: (1) observed evader states history, (2) self-observed states history, and (3) relative positions history between agents. These inputs are transformed using independent fully connected layers, aggregated through weighted averaging, and combined into a unified team representation. The final embedding is a fixed-dimensional vector, integrated into the actor network of MAPPO to enhance decision-making in adaptive teaming scenarios. The details of hyperparameters are listed in Table 5.

### v2

@@ 5 Conclusion
In this paper, we introduced AT-Drone, a novel multi-robot collaboration benchmark specifically designed to investigate adaptive teaming problems in multi-drone pursuit scenarios. AT-Drone uniquely integrates customizable simulation environments, real-world deployment pipelines leveraging Crazyflie drones and edge computing, a distributed training framework supporting diverse adaptive teaming algorithms, and standardized evaluation protocols. To facilitate comprehensive evaluation, AT-Drone provides four progressively challenging adaptive teaming environments and a collection of three distinct, unseen drone teammate zoos. Experimental results conducted with Crazyflie drones confirm AT-Drone’s effectiveness in driving advancements in adaptive teaming research. Furthermore, real-world drone experiments validate the benchmark’s practical feasibility and utility, demonstrating its relevance and applicability for realistic robotic operations.
6 Limitations
While AT-Drone successfully bridges simulation and real-world deployment, the current real-world system remains relatively simple, potentially limiting its ability to capture more intricate scenarios encountered in practical operations. Additionally, scalability constraints posed by physical hardware limitations—such as size, payload restrictions inherent to Crazyflie drones, and the current maximum of four pursuers and two evaders within a 3.6m × 5m area—may restrict larger-scale experimentation, making it challenging to study more complex pursuit tasks. Another critical limitation relates to perception; the current setup may not adequately handle complex perception tasks, potentially limiting the drones’ adaptability in visually challenging environments. Despite efforts to replicate realistic scenarios, simulation environments may not fully encapsulate all complexities and uncertainties present in real-world conditions. Future research should focus on addressing these limitations by enhancing system complexity, improving perception capabilities, exploring scalable drone hardware alternatives, enhancing simulation fidelity, and developing methods suitable for larger-scale, more complex scenarios.
Acknowledge
We sincerely thank Jianghong Wang for constructive feedback during the preparation of this manuscript.
Appendix A Environment Configurator
Appendix B Unseen Drone Zoo
@@
Traditional Method: VICSEK Drone. Based on the commonly used VICSEK algorithm [39, 2, 24], the VICSEK Drone adopts a bio-inspired approach to mimic swarm-like behaviours. It computes and updates a velocity vector directed towards the evader, optimising the tracking path based on the agent’s current environmental state. To avoid nearby obstacles or agents, the VICSEK Drone applies repulsive forces with varying magnitudes. While the calculated velocity vector includes both magnitude and orientation, only the orientation is implemented in our experiments, making it a scalable and practical teammate model for multi-drone coordination.
@@
Appendix C HOLA-Drone (V2) Algorithm
Building on the concept of preference hypergraphs [24], we use the preference hypergraph to represent the population and assess the coordination ability of each node. The preference hypergraph $\mathcal{P}\mathcal{G}$ is derived from the hypergraph $\mathcal{G}$, where each node has a direct outgoing hyperedge pointing to the teammates with whom it achieves the highest weight in $\mathcal{G}$. Formally, $\mathcal{P}\mathcal{G}$ is defined by the tuple $(\Pi,\mathcal{E}_{\mathcal{P}})$, where the node set $\Pi$ represents the strategies, and $\mathcal{E}_{\mathcal{P}}$ denotes the set of outgoing hyperedges. As shown in the right subfigure of Fig. 6, the dotted line highlights the outgoing edge. For instance, node $2$ has a single outgoing edge $(2,3,5,4)$ because it achieves the highest outcome, i.e., a weight of 45, with those teammates in $\mathcal{G}$, as depicted in the left subfigure.
Intuitively, a node in $\mathcal{P}\mathcal{G}$ with higher cooperative ability will have more incoming hyperedges, as other agents prefer collaborating with it to achieve the highest outcomes. Therefore, we extend the concept of preference centrality [14] to quantify the cooperative ability of each node. Specifically, for any node $i\in\Pi$, the preference centrality is defined as
$\eta_{\Pi}(i)=\frac{d_{\mathcal{P}\mathcal{G}}(i)}{d_{\mathcal{G}}(i)},$ (1)
Definition C.1 (Preference Optimal).
Figure 7: Overview of our proposed HOLA-Drone (V2) algorithm.
The max-min preference oracle updates the learner set by solving:
where the objective $\phi_{\mathcal{M}}(\cdot)$ is derived using the extended curry operator [40], originally designed for two-player games, and is expressed as:
$\mathcal{N}^{*}=\arg\max_{\mathcal{N}}\phi(\mathcal{N},\mathcal{M}^{*}).$
To achieve robust adaptability and dynamic coordination in multi-agent systems, we integrate the max-min preference oracle into an open-ended learning framework, referred to as the HOLA-Drone (V2) algorithm. The HOLA-Drone (V2) algorithm dynamically adjusts the training objective as the population evolves, enabling continuous improvement and effective coordination with unseen partners. Unlike conventional fixed-objective training, the HOLA-Drone (V2) approach iteratively expands the strategy population $\Pi$ and refines the learner set $\mathcal{N}$. At each generation $t$, the framework recalibrates the training objective $\phi$ based on new extended population $\Pi_{t}$ to account for the evolving interactions among agents within the population.
As shown in Fig. 6, HOLA-Drone (V2) algorithm consists of two key modules: the min-step solver and the max-step trainer. At each generation $t$, the updated learner set $\mathcal{N}_{t}$ from the previous generation $t-1$ is incorporated into the population $\Pi_{t-1}$, resulting in an expanded population $\Pi_{t}$.
Table 3: Implementation hyperparameters of NAHT -D algorithm.
@@ Appendix C HOLA-Drone (V2) Algorithm
Appendix D NAHT-D algorithm
NAHT-D extends the MAPPO algorithm [38] by incorporating an additional teammate modeling network $f$. This network generates team encoding vectors to represent the behavioral characteristics of unseen teammates, improving coordination in multi-drone pursuit. The modeling network $f$ processes three types of inputs: (1) observed evader states history, (2) self-observed states history, and (3) relative positions history between agents. These inputs are transformed using independent fully connected layers, aggregated through weighted averaging, and combined into a unified team representation. The final embedding is a fixed-dimensional vector, integrated into the actor network of MAPPO to enhance decision-making in adaptive teaming scenarios. The details of hyperparameters are listed in Table 3.
Figure 8: Case Study: This example demonstrates the capture strategy executed by NAHT-D learners and unseen drone partners from unseen zoo 3 in the superhard environment 4p3e5o. The red circles denote pursuers, and the black squares represent evaders. In this scenario, four pursuers collaboratively surround all three evader (1), two pursuers capture one of evaders while other two pursuers continuously tighten their formation (2), and rest of two evaders are then successfully captured one by one (3 & 4)
Appendix E Case Study
This case study demonstrates how the NAHT-D learners effectively coordinate with their unseen drone partners to execute a multi-stage capture strategy. Fig. 8 illustrates key frames from the scenario. In Frame 1, four pursuers initiate a collaborative approach, positioning themselves strategically to encircle all three evaders while maintaining an adaptive formation. In Frame 2, two pursuers successfully capture one of the evaders while the other two tighten their formation, preventing the remaining evaders from escaping. In Frames 3 & 4, the remaining two evaders are captured one by one as the pursuers continue refining their positioning and coordination, effectively closing all escape routes.